OpenStudy (anonymous):

Which ratio forms a proportion with 27/105 ? (A) 9/37 (B) 18/70 (C) 18/96 (D) 18/63 I thought the answer was 9/35 but strangely that's not one of the choices. So either me or the writers of this quiz has made an error.

OpenStudy (diyadiya):

You are right =D

OpenStudy (anonymous):

so the writers made a mistake with the test then?

OpenStudy (diyadiya):

Yeah :D

OpenStudy (anonymous):

Thanks Diyadiya

OpenStudy (diyadiya):

OH WAIT

OpenStudy (anonymous):

huh?

OpenStudy (diyadiya):

18/70

OpenStudy (diyadiya):

18/70 also equals to 9/35

OpenStudy (diyadiya):

So that is the Answer

OpenStudy (diyadiya):

Sorry for the confusion!!!

OpenStudy (anonymous):

lol how does 18/70 equal 9/35 can you briefly explain?

OpenStudy (diyadiya):

Both 18 and 70 are divisible by 2 so they can be simplified to 9/35 \[\Large \frac{\frac{18}{2}}{\frac{70}{2}}=\frac{9}{35}\]

OpenStudy (anonymous):

oh so basically if 9/5 was an option than it would be right but since it's not the answer is 18/70